### v2

# Words Matter: Leveraging Individual Text Embeddings for Code Generation in CLIP Test-Time Adaptation

Abstract
Vision-language foundation models, such as CLIP, have shown unprecedented zero-shot performance across a wide range of tasks. Nevertheless, these models may be unreliable under distributional shifts, as their performance is significantly degraded. In this work, we explore how to efficiently leverage class text information to mitigate these distribution drifts encountered by large pre-trained vision-language models (VLMs) during test-time inference. In particular, we propose to generate pseudo-labels for the test-time samples by exploiting generic class text embeddings as fixed centroids of a label assignment problem, which is efficiently solved with Optimal Transport. Furthermore, the proposed adaptation method (CLIP-OT) integrates a multiple template knowledge distillation approach, which replicates multi-view contrastive learning strategies in unsupervised representation learning but without incurring additional computational complexity. Extensive experiments on multiple popular test-time adaptation benchmarks presenting diverse complexity empirically show the superiority of CLIP-OT, achieving performance gains of up to $7\%$ over recent state-of-the-art methods, yet being computationally and memory efficient [^1] Code available at https://github.com/ShambhaviCodes/CLIPOT.
1 Introduction
Large pre-trained vision-language models (VLMs), such as CLIP [33] and ALIGN [19], have emerged as a new paradigm shift in machine learning, revealing promising zero-shot transferability. Nevertheless, if the model is exposed to domain drifts between the training and test samples, its performance can be largely degraded [44, 37].
While a straightforward solution to bridge this gap consists in fine-tuning the trained model using domain-specific labeled data [22, 13], this strategy brings several limitations in real-world scenarios, which may hamper their scalability and deployment. First, adapting to new domains requires collecting labeled samples drawn from each distinct distribution. This might be impractical for specific domains and further hinder the real-time adaptation of the trained model for each input test sample. Furthermore, fine-tuning the model may undermine its desirable zero-shot capabilities [40].
Test-Time Adaptation (TTA) presents a realistic and practical scenario of unsupervised domain adaptation, where a pre-trained model requires real-time adaptation to new data to address unknown distribution drifts without access to supervisory signals [39, 17, 45]. Nevertheless, despite the recent rise of VLMs, and the popularity of TTA in more traditional deep models, such as CNNs and ViTs, the study of TTA in large pre-trained vision-language models remains less explored. Initial approaches, such as TENT [39], which simply minimizes a Shannon entropy objective, have been adopted in CLIP test-time adaptation [36]. In contrast, more recent strategies resort to pseudo-labels on the inference batch to guide the model adaptation [31]. In particular, the core idea of these later methods is to minimize a classification loss, typically the standard cross-entropy, between the generated pseudo-labels and model predictions, which guides the network updates over multiple iterations. Nonetheless, while this strategy is common in semi-supervised learning, where labeled instances are available for all categories, applying it naively in unsupervised scenarios, such as test-time adaptation, can lead to degenerate solutions [17].
In light of the above limitation, we propose a novel approach for CLIP test-time adaptation based on the simple observation that this task can draw inspiration from clustering. Nevertheless, instead of jointly learning the cluster centroids during training, these can be replaced by the fixed class prototypes derived from the text representations. By doing this, the unlabeled samples in a batch will be assigned to the most similar class prototypes, guided by the text embeddings, which serve as a strong class guidance without requiring explicit sample-wise supervision. Furthermore, as our primary goal is to match distributions between different modalities, i.e., class text embeddings and features from the test images, traditional prototype-based clustering methods may lead to suboptimal performance, as they typically assume an unimodal distribution for each cluster. In contrast, we frame the label assignment problem as an Optimal Transport problem, as it can more naturally model multi-modal distributions [23], and is efficiently solved using the Sinkhorn algorithm [9].
Our key contributions can be summarized as follows:
  • We propose to cast the pseudo-labeling strategy in CLIP test-time adaptation as an Optimal Transport assignment, which leverages the class text information available in vision-language models in the form of fixed cluster centroids without requiring further annotations.
  • To solve the label assignment task, we resort to the Sinkhorn algorithm, as it can handle multi-modal distributions and compute label assignments efficiently.
  • We introduce a multi-template knowledge distillation approach, which leverages richer information derived from different text prompts to better guide the adaptation without incurring significant computational overhead.
  • Experiments across 240 scenarios demonstrate the superiority of CLIP-OT over recent state-of-the-art methods, all while avoiding additional computational complexity.
2 Related Work
Test-Time Adaptation.  TTA adapts a pre-trained model to a stream of incoming unlabeled target domain data, processed in batches during testing [39, 47, 8, 30]. Existing approaches for TTA in more traditional deep models can be roughly categorized into: i) normalization-based methods, which leverage the statistics of the test data to adjust the BatchNorm statistics of the model [27, 35]; ii) entropy-based approaches [39, 29, 12], where the model is adapted optimizing the Shannon entropy of the predictions; and iii) pseudo-label strategies [25], which employ the test-time generated labels for supervising the model. With the advent of VLMs, several works have attempted to accommodate some of these techniques to adapt large pre-trained foundation models, particularly CLIP [36, 26, 31, 14], which mainly differs on the different parameter groups updated during adaptation, i.e., prompt tuning [36, 26, 10] or layer-norm [14, 31] strategies. Test-Time Prompt Tuning (TPT) [36] fine-tunes only the input text prompts, which are optimized by minimizing the entropy of the model predictions. To achieve this, TPT generates multiple random augmented views for each test image, incurring additional costs due to performing multiple forward passes through the CLIP vision encoder for each image. Furthermore, TPT is a unimodal approach, which restricts a multi-modal model such as CLIP from fully exploiting its multi-modal capabilities for adaptation. Vision-Text-Space Ensemble (VTE) [10], uses an ensemble of different prompts as input to the CLIP text encoder, where both text and visual encoders remain frozen. Nonetheless, keeping the vision encoder frozen makes it difficult for the model to adapt to images with severe noise effectively. Most recently, CLIPArTT [14] and WATT [31] propose to fine-tune the normalization layers instead, resorting to pseudo-labels computed from image-to-image and text-to-text similarities.
Clustering for unsupervised representation learning.  Jointly learning how to adapt a set of learnable parameters of a deep network and identify the class assignments can be viewed as clustering and unsupervised representation learning. Thus, our work is closely related to recent literature on deep clustering-based approaches [1, 3, 4, 16, 42, 43, 18, 5]. In [3], $k$-means assignments are leveraged as pseudo-labels to learn visual representations, strategy that is later employed in [4] to pre-train standard supervised deep models. Nevertheless, applying naive $k$-means risks collapsing to only a few imbalanced clusters, making it ineffective. The label assignment task is framed as an instance of the Optimal Transport problem in [1], which stresses that the formulation presented in [3] is not principled, since potential degenerate solutions are avoided via specific implementation choices. Caron et al. [5] propose to enforce consistency between cluster assignments produced for different image views, or augmentations, which avoids expensive pairwise comparisons typically performed in contrastive learning. While we consider a similar formulation to some of these works, several vital differences exist. In particular, all these works, e.g., [4, 1, 3] are used in unsupervised representation learning, and thus the prototypes are learned from the data distribution. In contrast, our approach leverages the text representations to enhance the cluster assignments. Note that text embeddings are accessible for free in VLMs, not incurring additional supervision, as the test image label remains unknown. A natural benefit from this strategy is that we do not need to resort to additional augmentation or “multi-views” strategies [5], which introduce a computational burden. Indeed, we can leverage individual text templates as “augmented views”, enhancing the representation learning of the adapted model without incurring additional costs. In fact, these embeddings are computed only once offline. In contrast, image augmentation or “multi-views” strategies must create the additional images for each incoming data batch, which means they perform one forward pass per new augmentation at test time.
3 Background
3.1 Vision-language models
[…]
$p(y=k|\mathbf{x}_{i})=\frac{\exp{(\mathbf{z}_{i}^{\top}\mathbf{t}_{k}/\tau)}}{\sum_{j=1}^{K}\exp{(\mathbf{z}_{i}^{\top}\mathbf{t}_{j}/\tau)}},$ (1)
where $\tau$ is a temperature parameter, whose value is learned during pre-training [33], and $\mathbf{z}^{\top}\mathbf{t}$ is the dot product operator.
3.2 Test-Time Adaptation using CLIP
Problem setting.  We address the problem of adapting a pre-trained VLM at test time. In particular, given a model trained on the source domain $\mathcal{D}_{S}$, our goal is to adjust this model online to the new target domain $\mathcal{D}_{T}$, where only unlabeled test data $\{\mathbf{x}_{i}\}^{N}_{i=1}$ is available, and which is received as a stream of batches, from which predictions must be provided.
Vanilla entropy minimization.  Inspired by the semi-supervised learning literature, a straightforward solution to leverage the predicted probability in Eq. 1 in TTA would be to adapt the model parameters based on an entropy minimization objective, similar to TENT [39]:
$\mathcal{L}(p)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}p(y=k|\mathbf{x}_{i})\log p(y=k|\mathbf{x}_{i}),$ (2)
with $N$ denoting the size of the test set, or batch. However, relying on the Shannon entropy [39] in the fully unsupervised case poses a significant risk, as it may potentially result in a degenerate solution, i.e., Eq. (2) might be trivially minimized by assigning all data points to a single, arbitrary label.
4 Proposed solution
The proposed approach aims to overcome the limitations in Eq. 2 objective by finding reliable pseudo-labels in a batch of test-time samples without supervision. To achieve this, we propose CLIP-OT. This novel TTA method aims to better leverage the text embeddings generated by the frozen CLIP text encoder, serving as strong class descriptors to yield label assignments for test samples through Optimal Transport (OT). Note that our motivation drastically differs from the existing self-supervised representation learning literature that also resorts to pseudolabels [3], where cluster centroids are iteratively updated based on the prior label assignments. Below, we elaborate on the details of our method, which is illustrated in Figure 1.
Learning objective.  To overcome the limitation of naively minimizing Eq. 2 in the fully unlabeled scenario, we first encode the model predictions as posterior distributions $q(y|\mathbf{x}_{i})$, which results in:
$\mathcal{L}(p,q)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}q(y=k|\mathbf{x}_{i})\log p(y=k|\mathbf{x}_{i}).$ (3)
4.1 Online pseudolabels assignments
In the following, we detail how to encode robust posteriors into Eq. 3 by exploiting multi-modal capabilities of CLIP and the test-time batch data distribution. Considering the probability is obtained as in Eq. 1, the objective in Eq. 3 can be expressed as:
$\mathcal{L}(p,q)=-\frac{1}{N}\sum_{i=1}^{N}\left[\frac{1}{\tau}\mathbf{z}_{i}^{\top}\mathbf{T}\mathbf{q}_{i}-\log\sum_{k=1}^{K}\exp(\frac{\mathbf{z}_{i}^{\top}\mathbf{t}_{k}}{\tau})\right],$ (4)
where $\mathbf{T}=[\mathbf{t}_{1},...,\mathbf{t}_{K}]$ represents the matrix containing the class text prototypes, and the logarithmic term in the right side does not depend on $\mathbf{q}$. If we now express Eq. 4 in its matrix form over all the test images, $\mathbf{Z}$, and let $\mathbf{Q}$ be a matrix with columns $\mathbf{q}_{i}$, we have the following objective:
$\max_{\mathbf{Q}\in\mathcal{Q}}\;tr(\mathbf{Q}^{\top}\mathbf{T}^{\top}\mathbf{Z}),$ (5)
where, following [9], we enforce the matrix $\mathbf{Q}$ to be an element of the transportation polytope:
$\mathcal{Q}\mathrel{:}=\{\mathbf{Q}\in^{K\times N}_{+}\mid\mathbf{Q}\mathbf{1}_{N}=\frac{1}{K}\mathbf{1}_{K},\mathbf{Q}^{\top}\mathbf{1}_{K}=\frac{1}{N}\mathbf{1}_{N}\},$ (6)
with $\mathbf{1}_{K}$ and $\mathbf{1}_{N}$ denoting the vectors of ones in dimension $K$ and $N$, respectively. The constraints on $\mathcal{Q}$ enforce that on average each prototype is selected at least $\frac{N}{K}$ times, encouraging $\mathbf{Q}$ to be a matrix with uniform marginals in both rows and columns. Note that such constraints “break” the direct dependency between the original prediction $p$ in Eq. 2 and the posterior $q$ in Eq. 4, since the latter consider the whole batch distribution and its marginal properties for the assignment. This results in more robust supervisory signals, later demonstrated empirically (Figure 2).
Figure 1: CLIP-OT. Our method leverages Optimal Transport (with the Sinkhorn algorithm) to yield soft-codes $Q^{*}_{m}$ (Eq. 9). Then, it minimizes the pseudo cross-entropy between $Q^{*}_{m}$ and the CLIP predictions $P$ as unsupervised loss during test-time adaptation (Eq. 3). Note that, at each test batch, our model runs for $m$ iterations, each leveraging a different text template $m$ to obtain the pseudo-codes $Q^{*}_{m}$.
Codes optimization.  As the objective function in Eq. (5) is linear, and the constraints defining $\mathcal{Q}$ are also linear, this is a linear program. Nevertheless, directly optimizing the above learning objective might be time-consuming, particularly as the number of data points and classes increases. To address this issue and allow a faster optimization, we apply the Sinkhorn algorithm [9], which integrates an entropic constraint, enforcing a simple structure on the optimal regularized transport. Hence, the optimization problem becomes:
$\max_{\mathbf{Q}\in\mathcal{Q}}\;tr(\mathbf{Q}^{\top}\mathbf{T}^{\top}\mathbf{Z})+\varepsilon\mathcal{H}(\mathbf{Q}),$ (7)
where $\mathcal{H}(\mathbf{Q})$ denotes the Shannon entropy function $\mathcal{H}(\mathbf{Q})=-\sum_{ij}\mathbf{Q}_{ij}\log\mathbf{Q}_{ij}$ and $\varepsilon$ controls its weight. Furthermore, to align with TTA assumptions, we work on batches by restricting the transportation polytope to the current batch [5], in contrast to other works that employ the full dataset [1]. Thus, the dimensionality of the $\mathbf{Q}$ matrix becomes $K\times B_{s}$, where $B_{s}$ denotes the batch size. Now, the soft codes $\mathbf{Q}^{*}$ are the solution of the problem presented in (7) over the set $\mathcal{Q}$, which can be efficiently optimized with a few iterations as:
$\mathbf{Q}^{*}=\text{Diag}(\mathbf{u}^{(t)})\exp\left(\frac{\mathbf{T}^{\top}\mathbf{Z}}{\lambda}\right)\text{Diag}(\mathbf{v}^{(t)}),$ (8)
where $\mathbf{u}$ and $\mathbf{v}$ are renormalization vectors in K and B respectively, with $t$ denoting the iteration.
Optimizing codes for individual textual views.  Very recent literature [31] has pointed out the limitations of using the average text prototypes, and has instead suggested to resort to multiple category embeddings, each obtained through the different text templates in $\mathcal{T}$. Thus, we modify the solution presented in Eq. 8 to be specialized for specific class template [^2] The text templates are the same as in WATT [31] – see Appendix A., where the codes for each template $m$ are:
$\mathbf{Q}^{*}_{m}=\text{Diag}(\mathbf{u}^{(t)}_{m})\exp\left(\frac{\mathbf{T}^{\top}_{m}\mathbf{Z}}{\lambda}\right)\text{Diag}(\mathbf{v}^{(t)}_{m}).$ (9)
The renormalization vectors in our setting are computed trough a small number of matrix multiplications using the iterative Sinkhorn-Knopp algorithm, [20], where in each iteration $\mathbf{u}^{(t)}_{m}=\mathbf{u}_{m}/((\exp(\mathbf{T}_{m}^{\top}\mathbf{Z}/\lambda)\mathbf{v}^{t-1}_{m})$ and $\mathbf{v}^{(t)}_{m}=\mathbf{v}_{m}/((\exp(\mathbf{T}_{m}^{\top}\mathbf{Z}/\lambda)\mathbf{u}^{t}_{m})$, with $\mathbf{v}^{0}_{m}=\mathbf{1}$. Furthermore, $\mathbf{T}_{m}=[\mathbf{t}_{1m},...,\mathbf{t}_{Km}]$ denotes the matrix containing the class text prototypes obtained from the $m$-th template. Note that the computational burden introduced by OT is negligible, representing roughly $1\%$ of the total runtime for a batch, as discussed in Section 5.3.
4.2 Updating CLIP through textual codes
Distilling knowledge from multiple text prototypes.  To optimize the objective in Eq. 3, we employ the soft label assignments, as they yield superior performance compared to their hard counterpart in other problems [5]. Now, to distill the knowledge of the more diverse, and richer representations derived from multiple text predictions per category to the model, the cross-entropy in Eq. 3 is optimized $M$ times, each one using a different template $m$, which can be defined for each image and template as $\ell(\mathbf{p}_{i},\mathbf{q}^{*}_{im})=-\mathbf{q}^{*}_{im}\log\mathbf{p}_{i}$, where $\mathbf{p}_{i}$ is obtained with the average text template as in Eq. 1. Thus, for a given test image $i$ in a batch, after each update produced by the $m$-th text template, the visual encoder produces novel visual embeddings $\mathbf{z}_{i}$, which are then used to generate new predictions $\mathbf{p}_{i}$ for $i$.
Whole learning strategy.  The final proposed algorithm consists in updating the affine parameters in Layer Normalization, such as in [39, 31, 14], and learning a label assignment matrix $\mathbf{Q}_{m}$ per text template by solving the optimization problem presented in Eq. (7). This is done iteratively by alternating two steps across the batches and text templates: i) with the learnable parameters fixed, we compute the visual features $\mathbf{Z}\in^{d\times B_{s}}$ of the test batch samples, and find $\mathbf{Q}_{m}^{*}$ through Eq. (9) by iterating the updates on $\mathbf{u}_{m}$ and $\mathbf{v}_{m}$; and ii) given the current label assignments $\mathbf{Q}_{m}^{*}$, and the softmax predictions for the test samples obtained with each average class prototype $\mathbf{t}_{k}$, the set of learnable parameters is updated by minimizing Eq. (3) w.r.t. the layer norm parameters, where stochastic gradient descent is applied over the whole batch. This process is repeated $M$ times. Finally, the class predictions on the batch are inferred with the updated model. A view of the proposed method is detailed in Algorithm 1.
Thus, in a nutshell, we generate multiple pseudo-labels per sample, each derived from one text template. Note that pseudo-labeling is a common strategy in other fields, such as self-supervised [49] or semi-supervised [2, 46] learning. Thus, while $\mathbf{Q}$, or pseudo-labels in general, may contain errors, using these to refine the model is a standard practice, which typically yields consistent improvements. We refer the reader to Appendix G for a detailed discussion of the benefits of our approach compared to pseudo-labeling in other learning scenarios.
Interpretation.  Given a batch of $B_{s}$ visual embeddings $\mathbf{Z}=[\mathbf{z}_{1},...,\mathbf{z}_{B}]$, we aim at optimizing a function to map these embeddings into the class prototypes $\mathbf{T}=[\mathbf{t}_{1},...,\mathbf{t}_{K}]$. If we let $\mathbf{Q}$ denote the codes (or mapping), our goal is optimizing $\mathbf{Q}$ to maximize the similarity between the image features and the class text prototypes. While in fully-supervised, or semi-supervised regimes, labeled data is available to generate the class prototypes, in our Test-Time Adaptation scenario this becomes more challenging, as no label information about the test images is provided. Nevertheless, in the context of large vision-language models, the text can be leveraged as a form of supervision to create the prototypes $\mathbf{T}$, which represent a robust representation of each class. Furthermore, the entropic constraint term controls the smoothness of the mapping, promoting equally distributed predictions over all classes in each batch, which avoids degenerate solutions. Lastly, resorting to multiple templates individually, instead of averaging their embeddings, fully leverage the diverse information present in each text prompt.
Algorithm 1 CLIP-OT adaptation for one corruption.
1:input: test dataset $\mathcal{X}=\{\mathbf{x}_{n}\}$ (images), text templates $\mathcal{T}$ (class descriptions), visual and text encoders ($\boldsymbol{\theta}$, $\phi$). // Split $\mathcal{X}$ into $B$ batches of size $B_{s}$. // Compute $K\times M$ text prototypes ($\boldsymbol{\phi}(\mathcal{T}),\forall m,k)$.
2:$\mathbf{T}\in\mathbb{R}^{d\times K\times M}=\left[\mathbf{t}_{km}\right]_{k=1,\dots,K,m=1,\dots,M}$
3:for sampled minibatch $\{\mathbf{x}_{i}\}_{i=1}^{B_{s}}$ do // 1 - Adaptation
4:     for each template $m$ in a random permutation of $\{1,2,\dots,M\}$ do    // Step 1: Generation of pseudo-codes $Q^{*}_{m}$.
5:         $\mathbf{T}_{m}\in\mathbb{R}^{d\times K}$ // class text embeddings for $m$.
6:         $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features ($\boldsymbol{\theta}_{\text{LN}}^{(m-1)}$).
7:         Compute codes $\mathbf{Q}^{*}_{m}$ // Eq. (9).
8:         // Step 2: Distill knowledge to the model.
9:         $\mathbf{t}_{k}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{t}_{km}\quad(\mathbf{t}_{k}\in\mathbb{R}^{d})$ // average class text embeddings over all M templates.
10:         $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict: $\mathbf{t}_{k},\forall k$, Eq. (1).
11:         Min. cross-entropy with $\mathbf{P}$ and $\mathbf{Q}^{*}_{m}$ // Eq. (3).
12:         $\theta_{\text{LN}}^{(m-1)}\rightarrow\theta_{\text{LN}}^{(m)}$ // Update layer norm (LN) of $\boldsymbol{\theta}$.
13:     end for // 2 - Inference (for all images in the batch).
14:     $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features (with $\boldsymbol{\theta}_{\text{LN}}^{(m)}$).
15:     $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict with $\mathbf{t}_{k},\forall k$, Eq. (1).
16:end for
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours $\Delta$
CIFAR-10 88.74 91.69 $\pm$0.10 88.06 $\pm$0.06 90.04 $\pm$0.13 91.41 $\pm$0.17 91.05 $\pm$0.06 93.15 $\pm$0.10 +2.10 $\blacktriangle$
CIFAR-10.1 83.25 87.60 $\pm$0.45 81.80 $\pm$0.27 86.35 $\pm$0.27 87.78 $\pm$0.05 86.98 $\pm$0.31 88.37 $\pm$0.10 +1.39 $\blacktriangle$
CIFAR-10C Gaussian Noise 35.27 41.27 $\pm$0.27 33.90 $\pm$0.08 59.90 $\pm$0.36 61.89 $\pm$0.24 63.84 $\pm$0.24 64.85 $\pm$0.14 +1.01 $\blacktriangle$
Shot Noise 39.67 47.20 $\pm$0.23 38.20 $\pm$0.02 62.77 $\pm$0.07 63.52 $\pm$0.08 65.28 $\pm$0.21 67.34 $\pm$0.21 +2.06 $\blacktriangle$
Impulse Noise 42.61 48.58 $\pm$0.31 37.66 $\pm$0.20 56.02 $\pm$0.16 57.13 $\pm$0.02 58.64 $\pm$0.11 62.27 $\pm$0.43 +3.63 $\blacktriangle$
Defocus Blur 69.76 77.12 $\pm$0.16 67.83 $\pm$0.28 76.74 $\pm$0.05 78.86 $\pm$0.09 78.94 $\pm$0.12 82.09 $\pm$0.05 +3.15 $\blacktriangle$
Glass Blur 42.40 52.65 $\pm$0.30 38.81 $\pm$0.12 61.77 $\pm$0.16 62.88 $\pm$0.06 65.12 $\pm$0.07 68.07 $\pm$0.02 +2.95 $\blacktriangle$
Motion Blur 63.97 71.25 $\pm$0.09 63.39 $\pm$0.13 76.01 $\pm$0.19 76.85 $\pm$0.26 77.81 $\pm$0.14 81.30 $\pm$0.18 +3.49 $\blacktriangle$
Zoom Blur 69.83 76.20 $\pm$0.19 68.95 $\pm$0.16 77.40 $\pm$0.20 79.35 $\pm$0.04 79.32 $\pm$0.07 83.13 $\pm$0.24 +3.81 $\blacktriangle$
Snow 71.78 78.29 $\pm$0.20 70.16 $\pm$0.10 77.29 $\pm$0.16 79.44 $\pm$0.09 79.79 $\pm$0.06 83.71 $\pm$0.17 +3.92 $\blacktriangle$
Frost 72.86 79.84 $\pm$0.09 72.39 $\pm$0.22 79.20 $\pm$0.08 80.13 $\pm$0.10 80.54 $\pm$0.12 83.40 $\pm$0.20 +2.86 $\blacktriangle$
Fog 67.04 77.39 $\pm$0.01 64.31 $\pm$0.28 75.74 $\pm$0.14 77.68 $\pm$0.07 78.53 $\pm$0.22 82.56 $\pm$0.08 +4.03 $\blacktriangle$
Brightness 81.87 87.78 $\pm$0.03 81.30 $\pm$0.18 86.59 $\pm$0.16 87.10 $\pm$0.10 87.11 $\pm$0.11 89.90 $\pm$0.06 +2.79 $\blacktriangle$
Contrast 64.37 79.47 $\pm$0.11 62.26 $\pm$0.31 77.82 $\pm$0.14 80.04 $\pm$0.24 81.20 $\pm$0.22 84.86 $\pm$0.11 +3.66 $\blacktriangle$
Elastic Transform 60.83 70.00 $\pm$0.25 56.43 $\pm$0.27 70.20 $\pm$0.01 71.76 $\pm$0.10 72.66 $\pm$0.15 76.08 $\pm$0.30 +3.42 $\blacktriangle$
Pixelate 50.53 63.74 $\pm$0.18 42.80 $\pm$0.40 66.52 $\pm$0.13 69.28 $\pm$0.09 71.11 $\pm$0.13 76.25 $\pm$0.05 +5.14 $\blacktriangle$
JPEG Compression 55.48 62.64 $\pm$0.14 53.67 $\pm$0.25 63.51 $\pm$0.14 66.49 $\pm$0.14 67.36 $\pm$0.28 70.03 $\pm$0.37 +2.67 $\blacktriangle$
Average 59.22 67.56 56.80 71.17 72.83 73.82 77.06 +3.24 $\blacktriangle$
CIFAR-100 61.68 69.74 $\pm$0.16 63.78 $\pm$0.28 69.79 $\pm$0.04 70.38 $\pm$0.14 70.74 $\pm$0.20 71.68 $\pm$0.09 +0.94 $\blacktriangle$
CIFAR-100C Gaussian Noise 14.80 14.38 $\pm$0.14 14.03 $\pm$0.10 25.32 $\pm$0.14 31.28 $\pm$0.03 32.07 $\pm$0.23 33.43 $\pm$0.13 +1.36 $\blacktriangle$
Shot Noise 16.03 17.34 $\pm$0.27 15.25 $\pm$0.17 27.90 $\pm$0.05 33.44 $\pm$0.11 34.36 $\pm$0.11 35.60 $\pm$0.19 +1.24 $\blacktriangle$
Impulse Noise 13.85 10.03 $\pm$0.13 13.01 $\pm$0.13 25.62 $\pm$0.09 29.40 $\pm$0.11 30.33 $\pm$0.03 30.94 $\pm$0.41 +0.61 $\blacktriangle$
Defocus Blur 36.74 49.05 $\pm$0.07 37.60 $\pm$0.17 49.88 $\pm$0.23 52.32 $\pm$0.28 52.99 $\pm$0.16 53.87 $\pm$0.22 +0.88 $\blacktriangle$
Glass Blur 14.19 3.71 $\pm$0.07 16.41 $\pm$0.02 27.89 $\pm$0.03 31.20 $\pm$0.12 32.15 $\pm$0.30 35.26 $\pm$0.22 +3.11 $\blacktriangle$
Motion Blur 36.14 46.62 $\pm$0.27 37.52 $\pm$0.23 47.93 $\pm$0.14 49.72 $\pm$0.15 50.53 $\pm$0.12 52.77 $\pm$0.07 +2.24 $\blacktriangle$
Zoom Blur 40.24 51.84 $\pm$0.15 42.99 $\pm$0.11 52.70 $\pm$0.06 54.72 $\pm$0.04 55.30 $\pm$0.22 56.71 $\pm$0.17 +1.41 $\blacktriangle$
Snow 38.95 46.71 $\pm$0.21 42.35 $\pm$0.13 49.72 $\pm$0.01 51.79 $\pm$0.04 52.77 $\pm$0.15 54.30 $\pm$0.19 +1.53 $\blacktriangle$
Frost 40.56 44.90 $\pm$0.27 43.31 $\pm$0.14 49.63 $\pm$0.12 53.04 $\pm$0.08 53.79 $\pm$0.31 54.92 $\pm$0.07 +1.13 $\blacktriangle$
Fog 38.00 47.31 $\pm$0.04 38.81 $\pm$0.17 48.77 $\pm$0.04 50.78 $\pm$0.24 51.49 $\pm$0.21 53.57 $\pm$0.02 +2.08 $\blacktriangle$
Brightness 48.18 60.58 $\pm$0.18 50.23 $\pm$0.11 61.27 $\pm$0.08 62.65 $\pm$0.25 63.57 $\pm$0.21 64.43 $\pm$0.31 +0.86 $\blacktriangle$
Contrast 29.53 45.90 $\pm$0.11 28.09 $\pm$0.09 48.55 $\pm$0.24 51.34 $\pm$0.10 52.76 $\pm$0.27 55.01 $\pm$0.29 +2.25 $\blacktriangle$
Elastic Transform 26.33 33.09 $\pm$0.08 28.12 $\pm$0.15 37.45 $\pm$0.08 39.97 $\pm$0.06 40.90 $\pm$0.43 43.79 $\pm$0.27 +2.89 $\blacktriangle$
Pixelate 21.98 26.47 $\pm$0.09 20.43 $\pm$0.14 33.88 $\pm$0.14 39.59 $\pm$0.09 40.97 $\pm$0.16 44.51 $\pm$0.32 +3.54 $\blacktriangle$
JPEG Compression 25.91 29.89 $\pm$0.07 28.82 $\pm$0.09 36.07 $\pm$0.32 38.99 $\pm$0.16 39.59 $\pm$0.08 40.83 $\pm$0.14 +1.24 $\blacktriangle$
Average 29.43 35.19 30.46 41.51 44.68 45.57 47.33 +1.76 $\blacktriangle$
[…]
Link to unsupervised representation learning.  The proposed distilling strategy takes inspiration from unsupervised representation learning [7, 28, 5], where comparing random crops of the same image has the potential to yield stronger image representations. In particular, these strategies leverage these novel “views” of the same image as an image augmentation strategy, whose integration improves the semantic quality of the learned image representations. Nevertheless, i) this strategy substantially increases the computational and memory burden [7, 28], ii) specific composition of data augmentation operations is crucial for learning good representations [7], and iii) low-quality crops can alter the model performance [5]. In contrast, exploiting multiple text embeddings to guide the model adaptation process serves as a similar strategy, while also alleviating these shortcomings. More concretely, the text embeddings are low dimensional vectors, which can be computed off-line as they are image agnostic, considerably reducing computation requirements. Furthermore, in the absence of supervision, we argue that diverse text descriptions associated with a given class convey more informative representations than individual crops or “views” of an image to learn more generalizable features. While images provide valuable visual information, their parts may lack the contextual richness necessary for effective representation learning. This distinction suggests that leveraging a variety of textual prompts can indeed serve as a more promising alternative to contrastive learning based on multiple image crops while also being more efficient.
5 Experiments
5.1 Setting
Datasets.  All methods are benchmarked on standard TTA datasets known for their complexity and task diversity. These datasets include corrupted modifications from well-known datasets proposed in [15], such as CIFAR-10C, CIFAR-100C, or Tiny-ImageNet-C. Also, we employ its non-corrupted versions, i.e., CIFAR-10 and CIFAR-100 [21], CIFAR-10.1 [34], and Tiny-ImageNet [41]. Additionaly, we include Domainbed datasets (PACS [24], OfficeHome [38], VisDA-C [32]). More details can be found in Appendix B.
Baselines.  We resort to relevant CLIP-based TTA parameter-efficient adaptation approaches, i.e., modify a set of learnable parameters, including TENT [39], TPT [36], CLIPArTT [14], and WATT [31] (both versions WATT-S and WATT-P).
Architectures.  We employed the Vision Transformer architecture as the backbone for our experiments. The choice of ViT is driven by its superior performance on image classification tasks due to its attention mechanisms, which effectively capture long-range dependencies and spatial relationships within images. If the opposite is not stated, experiments are carried out using ViT-B/32, although we also provide further evaluations with larger ViT-backbones, e.g., ViT-B/16.
Implementation details.  Following the existing literature [31, 14], we used a batch size of 128 in our experiments and a learning rate of $10^{-4}$. The only additional hyper-parameters specific to our formulation are the entropic constraint weight, $\epsilon$ in Eq. 7, and the number of repetitions in the Sinkhorn algorithm. These are fixed to the same values across all experiments. $\epsilon$ is set to $\epsilon=0.7$, whereas the latter is fixed to $T=3$, which was enough for convergence, as in previous works [3]. Finally, we ran the adaptation experiments for each corruption/dataset three times under different random seeds and averaged the obtained accuracy accordingly.
5.2 Main results
We start our empirical validation by comparing the proposed approach’s performance to existing state-of-the-art CLIP-based TTA approaches across the different datasets.
Evaluation in the presence of natural or no domain shift.  Table 1 reports the performance of the different methods for CIFAR-10, CIFAR-10.1, and CIFAR-100, which introduce a natural or no domain shift. The proposed method consistently outperforms existing approaches, even when the domain shift is small, highlighting the usefulness of resorting to our approach over CLIP zero-shot predictions.
Impact of the performance under common corruptions.  Furthermore, in Table 1, we report a detailed analysis of the results obtained on CIFAR-10C and CIFAR-100C, which show the superiority of CLIP-OT in adapting CLIP in the presence of common corruptions. In particular, compared to the baseline CLIP, our model brings performance gains of nearly $18\%$ in CIFAR-10C and CIFAR-100C, without additional supervision. These performance gains are similar when compared to other popular TTA methods, e.g., TENT, or TPT, with differences ranging from $10\%$ to $21\%$. While this gap is reduced compared to very recent approaches, i.e., both versions of WATT and CLIPArTT, per-corruption classification accuracy differences are still significant, with CLIP-OT outperforming these two approaches by up to $5\%$ compared to WATT, i.e., $5.14\%$ in pixelate (CIFAR-10C) or $3.11\%$ in Glass Blur (CIFAR-100C), and nearly $11\%$ compared to CLIPArTT, e.g., $10.62\%$ in pixelate and $8.11\%$ in Gaussian Noise. As average, CLIP-OT improves the performance of very recent state-of-the-art approaches by a substantial difference of $3.24\%$ (CIFAR-10C), and $1.76\%$ in the more challenging scenario of CIFAR-100C, which contains $\times$10 number of classes.
What if the number of classes further increases?  In the following, we assess the performance of all methods on Tiny-ImageNet and Tiny-ImageNet-C benchmarks, which bring additional complexity due to the increased amount of categories present. From these results (Table 2), we can observe that the findings reported in previous sections hold, i.e., CLIP-OT significantly outperforming very recent state-of-the-art TTA approaches. More concretely, compared to WATT-S, the average improvement over all the corruptions is above $7.0\%$, with our method showing performance gains superior to $10\%$ in multiple corruptions, e.g., Shot Noise or Pixelate, and above $4.5\%$ in all remaining corruptions. Given the complexity and additional challenges that Tiny-ImageNet-C brings compared to CIFAR versions, as well as the non-negligible improvement provided by our method, we believe that these results showcase the clear superiority of CLIP-OT in more challenging scenarios.
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-S NeurIPS’24 CLIP-OT Ours
Tiny-ImageNet 58.29 57.72 58.90 59.85 61.35 63.69$\blacktriangle$
Gaussian Noise 7.08 8.01 9.29 14.44 13.02 21.40 $\blacktriangle$
Shot Noise 9.41 10.04 11.70 17.44 15.94 24.90 $\blacktriangle$
Impulse Noise 3.44 4.18 4.85 10.37 6.90 17.34 $\blacktriangle$
Defocus Blur 21.71 24.53 27.56 31.46 29.91 35.39 $\blacktriangle$
Glass Blur 9.12 10.09 11.03 15.84 14.01 21.16 $\blacktriangle$
Motion Blur 34.52 36.94 38.97 41.34 41.26 46.26 $\blacktriangle$
Zoom Blur 27.44 29.48 34.29 35.06 33.96 40.93 $\blacktriangle$
Snow 32.51 32.20 34.45 36.86 37.76 42.32 $\blacktriangle$
Frost 36.33 35.72 37.13 38.20 39.65 44.97 $\blacktriangle$
Fog 25.94 27.46 28.89 33.44 32.13 38.60 $\blacktriangle$
Brightness 40.15 39.79 43.31 46.43 46.93 53.10 $\blacktriangle$
Contrast 1.81 2.24 3.15 6.24 3.53 11.88 $\blacktriangle$
Elastic Transf. 30.40 31.92 33.88 33.89 35.01 40.73 $\blacktriangle$
Pixelate 22.78 24.79 27.70 34.85 31.55 41.84 $\blacktriangle$
JPEG Compr. 29.59 30.93 33.60 37.32 36.46 42.86 $\blacktriangle$
Mean 22.14 23.22 25.32 28.88 27.87 34.91 $\blacktriangle$
Table 2: Performance comparison on Tiny-ImageNet and Tiny-ImageNet-C corruption benchmarks using ViT-B/32 backbone.
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours $\Delta$
PACS 93.65 93.81 93.23 93.95 94.81 94.80 95.94 +1.14 $\blacktriangle$
OfficeHome 77.53 77.68 77.20 77.56 78.88 78.83 80.15 +1.32 $\blacktriangle$
VisDA-C (3D) 84.43 84.86 79.35 85.09 85.42 85.36 90.73 +5.37 $\blacktriangle$
VisDA-C (YT) 84.45 84.68 83.57 84.40 84.57 84.69 85.44 +0.75 $\blacktriangle$
Table 3: Performance comparison across PACS, OfficeHome, and VisDA-C datasets using ViT-B/32 backbone. $\Delta$ represents the difference between CLIP-OT and the State-of-the-art WATT-S [31]. Extended numerical values are presented in Appendix C.2, Table 9.
Performance under simulated and video shifts.  Table 3 (VisDA-C) reports the results of simulated (3D) and video (YT) shifts, which showcase the superiority of CLIP-OT compared to other methods. More concretely, the average difference compared to SoTA (e.g., WATT-S) is $3.03\%$, whereas per-dataset differences can go up to $11\%$, e.g., $11.38\%$ compared to TPT in 3D domain.
Texture and style shifts.  We now evaluate CLIP-OT on different distributional drifts, which may bring additional complexity. The results from these experiments (Table 3) show that the trend of superior performance is observed in almost all datasets (i.e., PACS and OfficeHome).
5.3 Ablation studies
Figure 2: Ablation on each component across multiple datasets. Different results come from adding each element of the proposed formulation. In green, performance differences compared to CLIP. The numerical values are provided in Appendix C.
Impact of each component.  In Figure 2, we empirically validate the benefit of each component in CLIP-OT design. Each configuration is detailed in Appendix C.1. First, we can observe that computing the code $Q^{*}_{m}$ from each template $m$ (Eq. (9)) and averaging those to obtain the final predictions consistently improves over zero-shot CLIP. It is noteworthy to highlight that the transportation polytope $\mathcal{Q}$ enforces additional explicit constraints on $\mathbf{Q}$, which ensure a well-defined probabilistic assignment and “break” the direct dependency between $\mathbf{Q}$ and $\mathbf{P}$. This explains the differences between direct CLIP predictions and this version of our approach, which directly leverages $\mathbf{Q}$ as predictions. Furthermore, while this performance is far from state-of-the-art approaches, this strategy is training-free, not requiring additional computational complexity due to gradient descent updates. Next, Average Template employs the pseudo-codes computed in the previous version to minimize the cross-entropy in Eq. (3). Furthermore, the prediction for the $i$-th image, $\mathbf{p}_{i}$, is computed with the average template for the class $k$, i.e., $\mathbf{t}_{k}$. This shows that leveraging the obtained pseudo-codes to refine the model leads to a performance improvement, which is consistent across all datasets. Last, we introduce the novel knowledge distillation strategy proposed in this work, where distinct templates are used to get multiple pseudo-codes and guide the model training, i.e., CLIP-OT. Its substantial performance gains are especially notable in the most challenging scenarios, e.g., CIFAR-100C.
On the impact of multiple templates.  We now assess how the performance evolves as the number of templates increases (Table 4). Concretely, we trained our model multiple times, varying the number of templates $N\in\{1,\dots,8\}$. For each $N$, the average accuracy over all template combinations was computed, leading to 256 experiments per dataset. Results from this experiment show that CLIP-OT performance consistently increases with the number of templates. This showcases how leveraging multiple templates individually can extract richer information to adapt CLIP at test time.
##\## Templates 1 2 3 4 5 6 7 8
CIFAR-10 91.59 92.16 92.46 92.69 92.88 93.04 93.14 93.14
CIFAR-100 68.36 69.46 70.08 70.56 70.90 71.15 71.38 71.34
Tiny-ImageNet 60.62 61.73 62.32 62.69 62.95 63.16 63.20 63.11
Table 4: Accuracy (%) of CLIP-OT with varying numbers of templates (see Appendix A) on natural domain shift benchmarks.
Do observations hold across backbones?  Table 5 reports the performance across different datasets when ViT-B/16 is employed as the visual encoder of CLIP. These results demonstrate that CLIP-OT is model agnostic, as its superiority is consistently maintained across different backbones.
Dataset CLIP TENT TPT CLIPArTT WATT-S CLIP-OT
CIFAR-10 89.25 92.75 89.90 92.61 91.97 94.69
CIFAR-10.1 84.00 88.52 83.75 88.72 88.10 89.25
CIFAR-10C 60.15 68.00 59.75 73.22 76.22 80.11
CIFAR-100 64.76 71.73 67.15 71.34 72.85 74.06
CIFAR-100C 32.01 37.90 33.73 40.08 48.95 51.24
TinyImageNet 58.12 63.59 63.87 64.73 66.05 66.34
TinyImageNet-C 20.92 29.78 26.96 32.90 31.66 37.69
Table 5: Quantitative results with ViT-B/16 as backbone for TTA across different datasets and baselines.
Computational analysis.  Figure 3 depicts the running time required for relevant baselines and the proposed CLIP-OT across several datasets, ordered in terms of increasing complexity, i.e., number of categories. These results expose that different TTA methods, particularly SoTA, substantially differ in total runtimes. In particular, recent CLIPArTT and WATT-S constantly increase the required runtime with the number of classes, driven by their iterative nature on the multiple template evaluations and weight averaging during inference. In contrast, CLIP-OT avoids this overhead by distilling this information during the adaptation stage and computing multiple text embeddings off-line only once. Moreover, the Sinkhorn algorithm—used for generating pseudo-codes is highly efficient, as its iterative updates converge faster on the low-dimensional similarity matrices and benefit from parallelizable computations, accounting for only nearly $1\%$ of the total runtime. This combination of speed and consistency establishes CLIP-OT as a scalable and effective method, providing runtimes from one to two orders of magnitude lower than CLIPArTT and WATT-S for the most complex dataset (i.e., Tiny-ImageNet), while significantly outperforming them from a discriminative standpoint.
Figure 3: Computational efficiency. The runtime of recent CLIP TTA methods is assessed, using a standard input batch of size 128. Runtime differences compared to our CLIP-OT highlighted in red.
6 Conclusions
We have presented a novel approach for TTA of VLMs. In particular, CLIP-OT generates pseudo-labels for test samples using multiple, generic class templates, optimizing a label assignment problem solved with Optimal Transport. This strategy yields state-of-the-art results compared to recent works in several benchmarks, yet being a significantly more efficient solution in terms of computational complexity.
Supplementary Material
Appendix A Templates Used
In the experimental setup, several predefined text templates from CLIP [11] were used to evaluate the proposed model’s adaptability and performance. These are designed to generalize across different contexts and image types, enabling a robust assessment of the model’s capabilities in handling diverse visual and textual representations. In Table 6, we provide the templates used in our experiments. Each template includes a placeholder , which is dynamically replaced with the class name during the generation of text prompts. Note that, for the sake of fairness with prior literature, the templates are the same as the ones employed in WATT [31].
Template
1: a photo of a {class $k$}
2: itap of a {class $k$}
3: a bad photo of the {class $k$}
4: a origami {class $k$}
5: a photo of the large {class $k$}
6: a {class $k$} in a video game
7: art of the {class $k$}
8: a photo of the small {class $k$}
Table 6: The different templates used during the experiments.
Appendix B Additional details on the datasets
CIFAR-10.1 introduces a natural shift from CIFAR-10, whereas CIFAR-10-C and CIFAR-100-C are augmented with 15 different corruptions across 5 severity levels (each containing 10,000 images), leading to 75 corruption scenarios commonly employed in domain shift problems. These datasets are critical for evaluating how well models can generalize to real-world variations and noise that are not present in the training set. Tiny-ImageNet is a downsized version of the original ImageNet dataset, providing a more accessible challenge with 200 classes. It is a useful benchmark for lightweight models and educational purposes, making it easier to implement and experiment with domain adaptation techniques. Tiny-ImageNet-C further extends Tiny-ImageNet by incorporating various common corruptions, similar to CIFAR-10-C and CIFAR-100-C. This dataset is instrumental for testing the robustness of models against a variety of distortions in a more manageable setting. To assess the performance on class imbalanced datasets, we employ several datasets from Domainbed (PACS, OfficeHome, VisDA-C) that are often utilized for benchmarking domain adaptation algorithms, and have been recently adopted in TTA scenarios [14]. Each dataset presents unique domain shift challenges and diverse visual categories, allowing for a comprehensive evaluation of a model’s ability to generalize across different environments.
Appendix C Additional details of the ablation studies
C.1 Configuration: Baselines
In this section, we detail the configurations used to assess the impact of each main component of our approach, which we refer to Training-free OT and CLIP-OT with Average Template. Note that, in short, Training-free OT motivates the use of the pseudo-codes $\mathbf{Q}^{*}$ over the CLIP baseline, whereas CLIP-OT with Average Template includes these pseudo-codes to fine-tune the model. Furthermore, we want to highlight that the later does not fully leverage multiple individual templates, which is introduced in our proposed CLIP-OT model. A more detailed description follows. Given a batch of test images, Training-free OT (Algorithm 2) computes pseudo-codes for each text template, which are later averaged to produce a final prediction. On the other hand, CLIP-OT with Average Template (Algorithm 3) utilizes the averaged pseudo-codes (obtained by Training-free OT) to refine the visual encoder. More concretely, at each batch, the average pseudo-codes $\tilde{\mathbf{Q}}$ supervise the predictions of the test images produced by the model. These predictions are obtained by resorting to Eq. (1), where the average class embedding $\mathbf{t}_{k}$ is used, together with the visual embeddings of the test images, $\mathbf{Z}$. Differences between the pseudo-codes and the predictions are minimized via a cross-entropy loss, whose gradients are used to update the layer norm parameters of the model. Then, once the model is updated, we can do the final inference for the test images.
C.2 Extended numerical values
We further substantiate the findings presented in Figure 2 by providing detailed results in Tables 7 and 8. The tables showcase a comprehensive performance comparison of the impact of the different components of our approach, which empirically motivate our model. These results are reported for different corruption benchmarks on the CIFAR-10, CIFAR-10.1, CIFAR-10C, CIFAR-100, and CIFAR-100C datasets, respectively, using a ViT-B/32 backbone. Also, Table 9 extends the results obtained on the different domains of PACS and OfficeHome, whose average results were reported in the main paper. We can observe that only in one out of the different datasets and domains, CLIP-OT underperforms previous SoTA, with a marginal performance gap, i.e., -0.19%.
Algorithm 2 Training-free OT.
1:input: test dataset $\mathcal{X}=\{\mathbf{x}_{n}\}$ (images), text templates $\mathcal{T}$ (class descriptions), visual and text encoders ($\boldsymbol{\theta}$, $\phi$). // Split $\mathcal{X}$ into $B$ batches of size $B_{s}$. // Compute $K\times M$ text prototypes ($\boldsymbol{\phi}(\mathcal{T}),\forall m,k)$.
2:$\mathbf{T}\in\mathbb{R}^{d\times K\times M}=\left[\mathbf{t}_{km}\right]_{k=1,\dots,K,m=1,\dots,M}$
3:for sampled minibatch $\{\mathbf{x}_{i}\}_{i=1}^{B_{s}}$ do
4:       for each template $m$ in a random permutation of $\{1,2,\dots,M\}$ do// Step 1: Generation of pseudo-codes $Q^{*}_{m}$.
5:             $\mathbf{T}_{m}\in\mathbb{R}^{d\times K}$ // class text embeddings for $m$.
6:             $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features ($\boldsymbol{\theta}_{\text{LN}}^{(m-1)}$).
7:             Compute pseudo-codes $\mathbf{Q}^{*}_{m}$ // Eq. (9).
8:       end for // Inference (for all images in the batch).
9:       $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict: $\mathbf{p}_{i}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{q}_{im}$.
10:end for
Algorithm 3 CLIP-OT with Average Template.
1:input: test dataset $\mathcal{X}=\{\mathbf{x}_{n}\}$ (images), text templates $\mathcal{T}$ (class descriptions), visual and text encoders ($\boldsymbol{\theta}$, $\phi$). // Split $\mathcal{X}$ into $B$ batches of size $B_{s}$. // Compute $K\times M$ text prototypes ($\boldsymbol{\phi}(\mathcal{T}),\forall m,k)$.
2:$\mathbf{T}\in\mathbb{R}^{d\times K\times M}=\left[\mathbf{t}_{km}\right]_{k=1,\dots,K,m=1,\dots,M}$
3:for sampled minibatch $\{\mathbf{x}_{i}\}_{i=1}^{B_{s}}$ do // 1 - Adaptation
4:       for each template $m$ in a random permutation of $\{1,2,\dots,M\}$ do// Step 1: Generation of pseudo-codes $Q^{*}_{m}$.
5:             $\mathbf{T}_{m}\in\mathbb{R}^{d\times K}$ // class text embeddings for $m$.
6:             $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features ($\boldsymbol{\theta}_{\text{LN}}^{(m-1)}$).
7:             Compute pseudo-codes $\mathbf{Q}^{*}_{m}$ // Eq. (9).
8:       end for // Step 2: Compute average pseudo-codes and refine the encoder.
9:       $\tilde{\mathbf{Q}}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{Q}^{*}_{m}$ // average pseudo-codes.
10:       $\mathbf{t}_{k}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{t}_{km}$ // average class text embeddings over all templates.
11:       $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict with $\mathbf{t}_{k},\forall k$, Eq. (1).
12:       Min. cross-entropy loss between $\mathbf{P}$ and $\tilde{\mathbf{Q}}$ // Eq. (3).
13:       $\theta_{\text{LN}}^{(t)}\rightarrow\theta_{\text{LN}}^{(t+1)}$ // update layer norm. // 2 - Inference (for all images in the batch).
14:       $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features (with $\boldsymbol{\theta}_{\text{LN}}^{(m)}$).
15:       $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict with $\mathbf{t}_{k},\forall k$, Eq. (1).
16:end for
Dataset CLIP Training-Free OT Avg. Template CLIP-OT
CIFAR-10 88.74 89.44 91.40 93.15
CIFAR-10.1 83.25 84.30 85.50 88.37
Gaussian Noise 35.27 46.68 54.68 64.85
Shot Noise 39.67 49.84 57.13 67.34
Impulse Noise 42.61 47.18 52.72 62.27
Defocus Blur 69.76 73.22 77.84 82.09
Glass Blur 42.40 50.09 56.88 68.07
Motion Blur 63.97 70.15 74.81 81.30
Zoom Blur 69.83 74.19 77.99 83.13
Snow 71.78 73.93 78.41 83.71
Frost 72.86 75.66 79.31 83.40
Fog 67.04 69.28 75.54 82.56
Brightness 81.87 82.90 86.87 89.90
Contrast 64.37 67.30 75.30 84.86
Elastic Transform 60.83 64.06 69.54 76.08
Pixelate 50.53 56.65 62.91 76.25
JPEG Compression 55.48 59.75 64.31 70.03
Mean 59.22 64.05 69.62 77.06
Table 7: Numerical Analysis for CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets as shown in Figure 2.
Corruption CLIP Training-Free OT Avg. Template CLIP-OT
CIFAR-100 61.68 61.55 68.26 71.68
Gaussian Noise 14.80 21.26 23.29 33.43
Shot Noise 16.03 22.98 25.15 35.60
Impulse Noise 13.85 22.27 22.14 30.94
Defocus Blur 36.74 41.40 46.84 53.87
Glass Blur 14.19 22.47 23.60 35.26
Motion Blur 36.14 40.71 44.41 52.77
Zoom Blur 40.24 45.14 49.97 56.71
Snow 38.95 44.14 46.73 54.30
Frost 40.56 45.03 47.85 54.92
Fog 38.00 40.00 45.45 53.57
Brightness 48.18 51.26 57.22 64.43
Contrast 29.53 33.69 41.16 55.01
Elastic Transform 26.33 33.02 34.85 43.79
Pixelate 21.98 27.72 30.13 44.51
JPEG Compression 25.91 31.69 33.80 40.83
Mean 29.43 34.85 38.17 47.33
Table 8: Numerical Analysis for CIFAR-100 and CIFAR-100-C datasets as shown in Figure 2.
Dataset Domain CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours $\Delta$
PACS Art 96.34 96.65 $\pm$0.05 95.52 $\pm$0.20 96.57 $\pm$0.09 96.31 $\pm$0.00 96.39 $\pm$0.00 96.54 $\pm$0.05 +0.15 $\blacktriangle$
Cartoon 96.08 96.22 $\pm$0.05 94.77 $\pm$0.20 96.00 $\pm$0.02 96.52 $\pm$0.02 96.62 $\pm$0.02 98.34 $\pm$0.13 +1.72 $\blacktriangle$
Photo 99.34 99.40 $\pm$0.00 99.42 $\pm$0.06 99.28 $\pm$0.00 99.48 $\pm$0.03 99.52 $\pm$0.00 99.86 $\pm$0.03 +0.34 $\blacktriangle$
Sketch 82.85 82.96 $\pm$0.12 83.22 $\pm$0.14 83.93 $\pm$0.14 86.92 $\pm$0.04 86.65 $\pm$0.12 89.01 $\pm$0.29 +2.36 $\blacktriangle$
Mean 93.65 93.81 93.23 93.95 94.81 94.80 95.94 +1.14 $\blacktriangle$
OfficeHome Art 73.75 74.03 $\pm$0.27 75.76 $\pm$0.27 73.84 $\pm$0.20 75.65 $\pm$0.27 75.76 $\pm$0.39 79.39 ±$\pm 0.31$ +3.63 $\blacktriangle$
Clipart 63.33 63.42 $\pm$0.04 63.08 $\pm$0.31 63.54 $\pm$0.06 66.23 $\pm$0.13 65.77 $\pm$0.11 66.21 $\pm$0.17 +0.44 $\blacktriangle$
Product 85.32 85.51 $\pm$0.08 84.07 $\pm$0.28 85.23 $\pm$0.16 85.41 $\pm$0.09 85.41 $\pm$0.01 86.81 $\pm$0.03 +1.40 $\blacktriangle$
[…]
Mean 77.53 77.68 77.20 77.56 78.88 78.83 80.15 +1.32 $\blacktriangle$
VisDA-C 3D (trainset) 84.43 84.86 $\pm$0.01 79.35 $\pm$0.04 85.09 $\pm$0.01 85.42 $\pm$0.03 85.36 $\pm$0.01 90.73 $\pm$0.65 +5.37 $\blacktriangle$
YT (valset) 84.45 84.68 $\pm$0.01 83.57 $\pm$0.04 84.40 $\pm$0.01 84.57 $\pm$0.00 84.69 $\pm$0.01 85.44 $\pm$0.65 +0.75 $\blacktriangle$
Mean 84.44 84.77 81.46 84.75 85.00 85.03 88.08 +3.05 $\blacktriangle$
[…]
Appendix D Comparison to OT-VP
The very recent OT-VP [48] has presented a solution integrating optimal transport for test-time adaptation. Nevertheless, it presents several fundamental differences with our work. First, OT-VP involves learning a universal visual prompt for the target domain, for which an optimal transport distance is optimized. Thus, optimal transport is used for a different task. Second, OT-VP is tailored to only visual models, not being capable of leveraging the valuable information found on the text modality. And third, the adaptation scenario strongly differs from our setting. In particular, OT-VP first fine-tunes the pre-trained model (pre-trained on ImageNet-1k) to a subdomain (e.g., PACS) with labeled data, to later adapt at test-time to the other subdomains (e.g., OfficeHome). While this strategy can be done in a single-source setting, OT-VP also evaluates the performance when multiple domains are used as the source (e.g., supervised adaptation on PACS), and only one left out for testing, referred to as Multi-Source in Table 10. In contrast, we follow the protocol of recent literature of CLIP test-time adaptation, where CLIP is directly exposed to the unsupervised test data points, without intermediate adaptation steps. To expand the extent of our empirical validation, we compare our approach to OT-VP [48] whose results are reported in Table 10. To ensure a rigorous comparison, we use the same visual encoder as the backbone used in OT-VP and other prior models, i.e., ViT-B/16.
Method PACS OfficeHome Avg.
First Setting : Single Source
[…]
Second Setting : Multi Source
[…]
Our setting: No specific source
CLIP-OT(Ours) 96.9+23.4/+9.2 83.4+15.3/+8.3 86.8+16.8/+5.6
Table 10: Performance comparison to [48] across PACS and OfficeHome datasets. Best results highlighted in bold, whereas value indicates the difference wrt OT-VP in both single (first value) and multi (second value) source scenarios.
Appendix E Differences wrt PLOT [6]
Note that despite the apparent similarities, our work presents fundamental differences w.r.t. Prompt Learning with Optimal Transport (PLOT) [6], which prevents from adding the later in the empirical evaluation. a) Target task. PLOT tackles few-shot adaptation, requiring labeled data to adapt no novel tasks, whereas our method studies test-time adaptation, where labels or test samples are not available. Thus PLOT cannot be compared in the TTA setting. b) Adaptation. Our approach updates the layer norm parameters of CLIP visual encoder, whereas PLOT optimizes the input text prompts (i.e., it falls into the Prompt Learning category). c) Use of OT. PLOT leverages OT between learnable prompts and a set of local visual features, whose transport plan is directly used as predicted probabilities. In contrast, we compute the transportation cost between the whole image features and fixed text embeddings, which is later used as a pseudo-supervision to distill knowledge during test-time adaptation.
Appendix F Additional Details on Epsilon ($\epsilon$)
In the main paper, we stated that $\epsilon$, the entropic constraint weight in Eq. 7 was set to $\epsilon=0.7$ based on preliminary experiments. However, here in the appendix, we explore the sensitivity of the model to different values of $\epsilon$ to assess its impact on accuracy under various corruptions and datasets. As shown in Table 11, we experimented with $\epsilon$ values of 0.3, 0.5, 0.7 and 0.9 across CIFAR-10C and CIFAR-100C datasets. Further, we show the values for CIFAR-10 and CIFAR-100 datasets along with averaged performance over corruptions in the Figure 4.
Corruption CIFAR-10 CIFAR-100
0.3 0.5 0.7 0.9 0.3 0.5 0.7 0.9
Original 91.15 93.01 93.15 92.81 62.26 70.95 71.62 71.01
Gaussian Noise 56.16 63.65 64.43 65.26 13.74 32.03 33.2 32.44
Shot Noise 57.53 66.29 66.27 67.4 15.62 34.69 35.33 35.46
Impulse Noise 54.27 61.22 61.84 62.42 19.44 29.94 30.99 30.36
Defocus Blur 76.27 81.33 82.1 82.49 39.16 52.77 54.16 53.32
Glass Blur 57.63 66.75 67.55 67.96 17.06 34.46 35.24 34.89
Motion Blur 71.99 80.11 81.31 80.9 36.59 51.33 52.67 51.62
Zoom Blur 76.24 82.54 83.11 83.46 42.58 55.53 56.93 56.13
Snow 78.59 83.39 82.99 83.56 40.91 53.54 54.06 54.12
Frost 78.92 82.67 82.93 83.41 40.98 54.1 54.93 54.37
Fog 75.75 81.64 81.9 82.47 40.76 52.69 53.6 53.12
Brightness 85.97 89.59 89.72 89.71 48.31 63.27 64.87 63.57
Contrast 70.38 83.55 84.53 84.6 32.09 53.61 55.39 54.41
Elastic Transform 69.17 75.42 75.67 76.09 27.78 42.69 43.78 43.01
Pixelate 61.31 74.57 76.51 76.08 22.79 42.64 44.96 43.18
JPEG Compression 66.02 69.42 70.46 70.03 30.16 40.16 41.02 40.67
Mean 70.46 76.14 76.75 77.06 31.20 46.23 47.41 46.71
Table 11: Mean Accuracy for CIFAR-10 and CIFAR-100 under Different Corruptions and Epsilon Values
We can observe that higher $\epsilon$ values (0.7 and 0.9) improved stability and performance due to stronger regularization introduced by higher entropic constraints, which prevents degenerate solutions during Sinkhorn normalization. While smaller $\epsilon$ values sometimes performed better under mild corruptions, they exhibited significant instability during Sinkhorn normalization. In several cases, the $\mathbf{Q}$ matrix became NaN, especially under high noise corruptions, rendering the adaptation process unusable which can be attributed to insufficient regularization, causing numerical instabilities during iterative updates. The chosen value ($\epsilon=0.7$) strikes a balance between stability and performance, yielding robust results across both clean and corrupted datasets.
Figure 4: Ablation on epsilon values across multiple datasets.
Appendix G Detailed benefits of our multi-template distillation approach.
The proposed multi-template distillation strategy resembles to the unsupervised approach employed to leverage pseudo-labels in semi-supervised learning, e.g., [2]. In this learning scenario, unlabeled samples are augmented, whose predictions are averaged to obtain a pseudo-label as $\mathbf{q}_{u}=\frac{1}{M}\sum_{m}f_{\theta}(y|\bar{\mathbf{x}}_{u,m})$, where $\bar{\mathbf{x}}_{u,m}$ is the $m$-th augmented version of a given unlabeled image $\mathbf{x}_{u}$. Once the average pseudo-label is obtained, we can minimize the pseudo cross-entropy as $-\sum_{k}\mathbf{q}_{u}\log f_{\theta}(y|\bar{\mathbf{x}}_{u,m})$ [^3] For simplicity, we consider only one unlabeled image.. In this setting, the predictions from multiple augmented images are guided by the average pseudo-label. Once the loss is computed over all the predictions from the augmented images, the gradient is back-propagated and the network parameters updated. Limitations of this method: First, this approach forces all augmentations to align with a single static target $\mathbf{q}$, potentially causing the model to converge to a suboptimal solution because it cannot make incremental adjustments based on each individual augmentation. Second, it treats all augmentations as being equally aligned with a single target $\mathbf{q}$, which could limit the model’s ability to handle diverse augmentations, especially if some augmentations are very different from others. And last, this strategy applies a global update that may force the model too quickly into a solution where all augmentations have the same prediction target $\mathbf{q}$, potentially making it harder for the model to learn robust features. To overcome these limitations in the TTA scenario, where no single supervision is available, the proposed method takes an opposite direction to these well-established practices. More concretely, our approach minimizes the following pseudo cross-entropy loss at each step $m$: $-\mathbf{q}_{m}\log\mathbf{p}$, where $\mathbf{q}_{m}$ is the pseudo-label at step $m$ (derived from the $m$-th text prompt, which is equivalent to one augmentation), and $\mathbf{p}$ is the mean prediction considering all the $M$ text prompts (similar to averaging over $M$ augmentations). At step $m$, the gradient is backpropagated, the network parameters updated. Then, we can recompute the pseudo-labels at step $m+1$, as well as the new mean prediction $\mathbf{p}$ (which will be now different due to the parameters update). Advantages: First, by updating the model after each step and recomputing $\mathbf{p}$ as the mean of the model’s outputs, our approach encourages a more fine-grained, iterative refinement of predictions, potentially helping the model to converge more smoothly. Indeed, the model can adjust its predictions for each template (i.e., augmentation) before being evaluated again in the next step. Furthermore, this step-by-step refinement process allows the model to better adapt to the diversity in the augmentations and learn a more generalizable representation.
Appendix H Discussion on the Two Hyperparameters of WATT and Motivation for CLIP-OT
WATT [31] introduces two key hyperparameters: $L$ (number of adaptation iterations for each text embedding) and $M$ (number of repetitions of the weight averaging process). While these hyperparameters play an essential role in improving WATT’s adaptation capabilities, they introduce significant computational and scalability challenges, particularly as the number of classes grows. The iterative nature of $L$ requires multiple forward and backward passes for every template, leading to substantial runtime overhead, especially for datasets with many classes like Tiny-ImageNet. Additionally, higher $L$ values, while effective for complex corruptions, risk overfitting to noisy pseudo-labels, while lower $L$ values may lead to under-adaptation. Similarly, $M$ stabilizes adaptation through repeated weight averaging, but its effectiveness diminishes beyond a certain point, with larger values providing negligible performance improvements while exponentially increasing runtime. This dependence on repeated updates for each class and template becomes especially prohibitive for datasets with a large number of classes, as demonstrated in Figure 3, where runtime increases from CIFAR-10 (10 classes) to Tiny-ImageNet (200 classes).
These limitations motivate our proposed approach, CLIP-OT, which simplifies and optimizes the adaptation process, achieving robust performance without the drawbacks associated with $L$ and $M$. CLIP-OT eliminates these limitations by leveraging the Sinkhorn algorithm for single-pass optimization and precomputing averaged class embeddings, removing the need for iterative updates and weight averaging. This not only drastically reduces computational overhead but also ensures scalable, robust adaptation across diverse corruption levels and datasets, offering a practical alternative to WATT’s hyperparameter-dependent framework.
Appendix I Further results on other backbones
In this section, we present additional experimental results leveraging different backbone architectures to evaluate the robustness and generalizability of our proposed method. Specifically, we analyze the performance of CLIP models with ViT-B/16 and ViT-L/14 backbones across various datasets, including CIFAR, Tiny-ImageNet and their corruptions. These experiments validate our approach across diverse settings and highlight its adaptability to challenging domain shifts.
Table 12 details the performance of CLIP-OT with the ViT-B/16 backbone on CIFAR-10, CIFAR-10.1, and CIFAR-10-C datasets. Our proposed method demonstrates consistent improvements across both clean and corrupted data, outperforming baseline methods such as CLIP, TENT, and TPT. Similarly for ViT-L/14 backbone on these datasets, we obtain similar gains as illustrated in Table 13. Table 14 extends the evaluation to CIFAR-100 and CIFAR-100-C, demonstrating the effectiveness of our method on more complex datasets with larger label spaces. Using CLIP with the ViT-B/16 backbone, our approach achieves the highest average accuracy, with substantial improvements under corruptions such as Glass Blur ($+5.30\%$) and Contrast ($+6.23\%$). We also evaluate these datasets with backbone ViT/L-14 in Table 15.
For Tiny-ImageNet and its corrupted version, Tiny-ImageNet-C, Table 16 and Table 17 provides a detailed breakdown. On datasets with higher complexity and larger label spaces, such as Tiny-ImageNet and Tiny-ImageNet-C (Table 16), CLIP-OT exhibits substantial gains, consistent with the main paper’s observations. Specifically, our method improves over WATT-S by over 6% on average across Tiny-ImageNet-C corruptions, showcasing its scalability to more challenging scenarios.
I.1 Tiny-Imagenet Experiments
We report the mean and standard deviation over Tiny-Imagenet experiments (Table 2, main paper) in Table 18.
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours
CIFAR-10 89.25 92.75 $\pm$0.17 89.80 $\pm$0.05 92.61 $\pm$0.05 92.31 $\pm$0.10 91.97 $\pm$0.03 94.69 $\pm$0.18
CIFAR-10.1 84.00 88.52 $\pm$0.33 83.75 $\pm$0.21 88.72 $\pm$0.33 87.90 $\pm$0.11 88.10 $\pm$0.08 89.25 $\pm$0.08
Gaussian Noise 37.75 31.04 $\pm$0.38 35.35 $\pm$0.15 60.89 $\pm$0.26 63.10 $\pm$0.12 65.57 $\pm$0.22 68.50 $\pm$0.19
Shot Noise 41.10 40.54 $\pm$0.41 41.03 $\pm$0.19 65.19 $\pm$0.21 66.31 $\pm$0.10 68.67 $\pm$0.37 71.24 $\pm$0.22
Impulse Noise 51.71 58.03 $\pm$0.16 54.86 $\pm$0.07 67.55 $\pm$0.09 69.62 $\pm$0.12 70.39 $\pm$0.11 76.97 $\pm$0.17
Defocus Blur 70.07 77.57 $\pm$0.03 70.29 $\pm$0.02 78.92 $\pm$0.12 79.64 $\pm$0.08 79.90 $\pm$0.07 83.33 $\pm$0.09
Glass Blur 42.24 47.16 $\pm$0.05 37.86 $\pm$0.17 57.18 $\pm$0.20 58.98 $\pm$0.12 61.62 $\pm$0.21 68.33 $\pm$0.20
Motion Blur 65.81 76.16 $\pm$0.05 67.43 $\pm$0.11 76.59 $\pm$0.06 78.32 $\pm$0.16 79.02 $\pm$0.07 83.13 $\pm$0.10
Zoom Blur 72.50 79.64 $\pm$0.12 72.91 $\pm$0.02 79.62 $\pm$0.11 80.67 $\pm$0.07 81.10 $\pm$0.08 85.35 $\pm$0.11
Snow 73.23 81.68 $\pm$0.03 72.98 $\pm$0.32 81.13 $\pm$0.29 81.99 $\pm$0.10 82.54 $\pm$0.18 86.65 $\pm$0.12
Frost 76.52 83.22 $\pm$0.05 75.87 $\pm$0.16 81.24 $\pm$0.08 83.41 $\pm$0.16 83.46 $\pm$0.15 86.37 $\pm$0.09
Fog 68.35 80.78 $\pm$0.15 69.13 $\pm$0.27 78.47 $\pm$0.19 81.36 $\pm$0.12 81.88 $\pm$0.12 85.60 $\pm$0.24
Brightness 83.36 89.85 $\pm$0.11 83.67 $\pm$0.14 88.66 $\pm$0.15 89.06 $\pm$0.05 89.10 $\pm$0.14 92.64 $\pm$0.11
Contrast 61.90 79.24 $\pm$0.19 62.16 $\pm$0.06 75.15 $\pm$0.07 81.57 $\pm$0.23 83.79 $\pm$0.12 86.92 $\pm$0.18
Elastic Transform 53.16 62.54 $\pm$0.08 51.26 $\pm$0.23 69.49 $\pm$0.08 69.14 $\pm$0.09 70.93 $\pm$0.20 75.18 $\pm$0.14
Pixelate 48.48 67.08 $\pm$0.24 44.65 $\pm$0.21 71.80 $\pm$0.16 73.38 $\pm$0.29 75.67 $\pm$0.32 78.73 $\pm$0.30
JPEG Compression 56.05 65.42 $\pm$0.05 56.73 $\pm$0.07 66.42 $\pm$0.25 69.02 $\pm$0.10 69.65 $\pm$0.23 72.65 $\pm$0.18
Mean 60.15 68.00 59.75 73.22 75.04 76.22 80.11
[…]
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours
CIFAR-10 95.36 96.13 $\pm$0.06 95.18 $\pm$0.02 95.16 $\pm$0.03 95.91 $\pm$0.10 95.71 $\pm$0.03 97.61 $\pm$0.03
CIFAR-10.1 91.20 92.22 $\pm$0.25 91.32 $\pm$0.12 91.02 $\pm$0.02 92.97 $\pm$0.13 92.10 $\pm$0.33 93.45 $\pm$0.30
Gaussian Noise 64.64 68.87 $\pm$0.20 64.44 $\pm$0.11 70.04 $\pm$0.31 72.81 $\pm$0.09 72.73 $\pm$0.03 80.47 $\pm$0.26
Shot Noise 67.82 71.95 $\pm$0.06 66.81 $\pm$0.19 71.44 $\pm$0.16 74.45 $\pm$0.16 74.60 $\pm$0.03 82.64 $\pm$0.29
Impulse Noise 78.21 80.22 $\pm$0.19 76.46 $\pm$0.17 79.42 $\pm$0.15 81.36 $\pm$0.09 80.95 $\pm$0.15 87.80 $\pm$0.15
Defocus Blur 80.73 83.10 $\pm$0.03 79.01 $\pm$0.23 81.75 $\pm$0.19 83.20 $\pm$0.10 83.15 $\pm$0.18 90.33 $\pm$0.12
Glass Blur 50.29 57.12 $\pm$0.07 49.64 $\pm$0.23 58.13 $\pm$0.23 61.51 $\pm$0.07 62.35 $\pm$0.15 77.72 $\pm$0.31
Motion Blur 80.75 82.69 $\pm$0.11 78.85 $\pm$0.04 80.76 $\pm$0.12 82.60 $\pm$0.13 82.61 $\pm$0.12 88.93 $\pm$0.17
Zoom Blur 82.75 84.91 $\pm$0.08 82.32 $\pm$0.13 83.39 $\pm$0.05 85.76 $\pm$0.06 85.44 $\pm$0.13 91.69 $\pm$0.13
Snow 83.01 85.99 $\pm$0.11 82.69 $\pm$0.10 84.48 $\pm$0.07 84.91 $\pm$0.13 85.61 $\pm$0.15 91.56 $\pm$0.11
Frost 84.90 87.15 $\pm$0.12 84.63 $\pm$0.08 85.21 $\pm$0.06 87.17 $\pm$0.13 86.88 $\pm$0.04 91.43 $\pm$0.12
Fog 78.44 81.30 $\pm$0.07 77.56 $\pm$0.17 79.27 $\pm$0.07 81.80 $\pm$0.11 81.79 $\pm$0.09 90.82 $\pm$0.11
Brightness 91.67 93.07 $\pm$0.04 90.94 $\pm$0.04 91.87 $\pm$0.09 92.78 $\pm$0.05 92.59 $\pm$0.16 95.86 $\pm$0.00
Contrast 84.20 87.93 $\pm$0.04 82.88 $\pm$0.09 86.19 $\pm$0.06 87.54 $\pm$0.12 87.38 $\pm$0.02 94.53 $\pm$0.08
Elastic Transform 65.45 69.96 $\pm$0.12 64.81 $\pm$0.14 67.43 $\pm$0.24 71.19 $\pm$0.07 71.25 $\pm$0.09 82.16 $\pm$0.23
Pixelate 75.10 77.89 $\pm$0.05 72.92 $\pm$0.12 77.11 $\pm$0.10 77.88 $\pm$0.13 77.67 $\pm$0.16 87.13 $\pm$0.16
JPEG Compression 72.58 75.49 $\pm$0.07 71.18 $\pm$0.19 74.46 $\pm$0.11 75.88 $\pm$0.16 75.84 $\pm$0.18 82.60 $\pm$0.18
Mean 76.04 79.18 75.01 78.06 80.05 80.06 86.35
[…]
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours
CIFAR-100 64.76 71.73 $\pm$0.14 67.15 $\pm$0.23 71.34 $\pm$0.07 72.98 $\pm$0.07 72.85 $\pm$0.15 74.07 $\pm$0.20
Gaussian Noise 15.88 12.28 $\pm$0.20 15.43 $\pm$0.03 19.01 $\pm$0.24 34.23 $\pm$0.03 35.95 $\pm$0.27 38.80 $\pm$0.51
Shot Noise 17.49 15.07 $\pm$0.21 16.88 $\pm$0.07 20.27 $\pm$0.21 36.68 $\pm$0.10 37.96 $\pm$0.15 40.89 $\pm$0.10
Impulse Noise 21.43 13.13 $\pm$0.16 22.12 $\pm$0.15 17.66 $\pm$0.10 43.17 $\pm$0.35 44.62 $\pm$0.20 46.39 $\pm$0.17
Defocus Blur 40.10 50.35 $\pm$0.03 41.08 $\pm$0.22 49.86 $\pm$0.13 53.13 $\pm$0.12 53.80 $\pm$0.12 55.50 $\pm$0.24
Glass Blur 13.48 04.84 $\pm$0.14 18.43 $\pm$0.15 18.34 $\pm$0.31 32.53 $\pm$0.03 33.39 $\pm$0.11 37.89 $\pm$0.68
Motion Blur 39.82 49.85 $\pm$0.37 40.85 $\pm$0.26 50.00 $\pm$0.09 51.63 $\pm$0.06 52.72 $\pm$0.30 55.10 $\pm$0.28
Zoom Blur 45.45 54.76 $\pm$0.04 46.77 $\pm$0.06 54.13 $\pm$0.08 56.81 $\pm$0.11 57.51 $\pm$0.09 58.77 $\pm$0.33
Snow 42.77 52.38 $\pm$0.18 47.24 $\pm$0.18 52.80 $\pm$0.27 56.04 $\pm$0.06 56.73 $\pm$0.27 57.84 $\pm$0.13
Frost 45.39 51.66 $\pm$0.04 48.61 $\pm$0.14 49.56 $\pm$0.08 56.00 $\pm$0.11 56.48 $\pm$0.34 58.09 $\pm$0.04
Fog 38.98 50.74 $\pm$0.14 39.92 $\pm$0.16 49.92 $\pm$0.11 52.88 $\pm$0.33 53.83 $\pm$0.19 55.68 $\pm$0.38
Brightness 52.55 64.26 $\pm$0.09 55.83 $\pm$0.10 63.76 $\pm$0.13 65.58 $\pm$0.07 66.67 $\pm$0.19 67.76 $\pm$0.21
Contrast 33.32 48.69 $\pm$0.08 33.13 $\pm$0.16 47.86 $\pm$0.02 52.90 $\pm$0.06 55.06 $\pm$0.15 58.29 $\pm$0.51
Elastic Transform 24.39 33.56 $\pm$0.28 27.36 $\pm$0.10 32.93 $\pm$0.23 39.82 $\pm$0.21 40.37 $\pm$0.26 44.32 $\pm$0.18
Pixelate 21.89 36.20 $\pm$0.28 21.26 $\pm$0.10 39.49 $\pm$0.21 45.10 $\pm$0.06 47.02 $\pm$0.04 49.46 $\pm$0.17
JPEG Compression 27.21 30.80 $\pm$0.05 30.97 $\pm$0.10 35.56 $\pm$0.23 41.43 $\pm$0.18 42.13 $\pm$0.21 43.87 $\pm$0.23
Mean 32.01 37.90 33.73 40.08 47.86 48.95 51.24
[…]
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours
CIFAR-100 73.28 78.03 $\pm$0.08 76.85 $\pm$0.06 79.42 $\pm$0.08 79.33 $\pm$0.05 78.85 $\pm$0.19 81.38 $\pm$0.15
Gaussian Noise 30.55 36.93 $\pm$0.03 36.10 $\pm$0.11 41.46 $\pm$0.15 43.99 $\pm$0.13 44.13 $\pm$0.11 52.26 $\pm$0.18
Shot Noise 34.58 40.96 $\pm$0.16 38.23 $\pm$0.13 44.27 $\pm$0.09 46.28 $\pm$0.22 46.63 $\pm$0.17 54.59 $\pm$0.26
Impulse Noise 44.89 49.09 $\pm$0.14 49.69 $\pm$0.21 51.44 $\pm$0.23 56.15 $\pm$0.04 56.26 $\pm$0.22 63.02 $\pm$0.12
Defocus Blur 48.88 55.23 $\pm$0.07 50.43 $\pm$0.19 56.55 $\pm$0.22 57.46 $\pm$0.01 57.66 $\pm$0.42 65.20 $\pm$0.17
Glass Blur 23.46 27.02 $\pm$0.23 24.35 $\pm$0.22 30.47 $\pm$0.14 32.54 $\pm$0.12 33.54 $\pm$0.16 48.19 $\pm$0.17
Motion Blur 50.83 56.03 $\pm$0.20 51.94 $\pm$0.04 56.98 $\pm$0.18 58.22 $\pm$0.10 57.81 $\pm$0.05 65.09 $\pm$0.17
Zoom Blur 56.02 61.19 $\pm$0.10 56.96 $\pm$0.16 62.56 $\pm$0.04 62.94 $\pm$0.02 62.74 $\pm$0.06 68.92 $\pm$0.15
Snow 49.03 55.60 $\pm$0.09 54.89 $\pm$0.11 58.81 $\pm$0.11 60.68 $\pm$0.06 61.04 $\pm$0.27 66.81 $\pm$0.15
Frost 53.27 58.21 $\pm$0.15 58.15 $\pm$0.33 60.38 $\pm$0.23 62.34 $\pm$0.14 62.76 $\pm$0.22 67.71 $\pm$0.29
Fog 48.51 53.37 $\pm$0.25 49.26 $\pm$0.13 54.38 $\pm$0.04 54.71 $\pm$0.31 54.70 $\pm$0.13 65.54 $\pm$0.23
Brightness 60.53 67.34 $\pm$0.17 66.60 $\pm$0.10 69.63 $\pm$0.14 71.52 $\pm$0.07 71.60 $\pm$0.09 76.25 $\pm$0.15
Contrast 50.24 59.91 $\pm$0.13 53.64 $\pm$0.24 63.39 $\pm$0.13 62.77 $\pm$0.22 63.95 $\pm$0.04 72.34 $\pm$0.27
Elastic Transform 35.07 38.49 $\pm$0.12 35.72 $\pm$0.09 39.57 $\pm$0.39 41.28 $\pm$0.25 41.27 $\pm$0.15 52.72 $\pm$0.24
Pixelate 43.86 48.37 $\pm$0.17 44.32 $\pm$0.10 50.45 $\pm$0.16 51.15 $\pm$0.15 51.22 $\pm$0.13 59.55 $\pm$0.06
JPEG Compression 39.11 44.42 $\pm$0.09 43.44 $\pm$0.11 47.45 $\pm$0.14 49.40 $\pm$0.17 49.78 $\pm$0.18 54.98 $\pm$0.20
Mean 44.59 50.14 47.58 52.52 54.10 54.34 69.48
[…]
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-S NeurIPS’24 CLIP-OT Ours
Tiny-ImageNet 58.12 63.59 63.87 $\pm$0.09 64.73 $\pm$0.11 66.05 $\pm$0.03 66.34 $\pm$0.07
Gaussian Noise 4.80 13.44 $\pm$0.21 10.53 $\pm$0.25 19.25 $\pm$0.18 16.78 $\pm$0.09 24.88 $\pm$0.28
Shot Noise 6.44 17.46 $\pm$0.20 12.90 $\pm$0.45 21.80 $\pm$0.11 19.27 $\pm$0.13 27.99 $\pm$0.32
Impulse Noise 4.36 9.09 $\pm$0.14 6.57 $\pm$0.20 13.46 $\pm$0.01 10.42 $\pm$0.07 20.69 $\pm$0.18
Defocus Blur 22.23 32.33 $\pm$0.09 29.16 $\pm$0.32 33.49 $\pm$0.29 $3$$\pm$0.03 37.27 $\pm$0.16
Glass Blur 7.50 12.47 $\pm$0.14 10.88 $\pm$0.18 18.33 $\pm$0.09 16.26 $\pm$0.08 22.72 $\pm$0.14
Motion Blur 33.46 44.17 $\pm$0.06 39.91 $\pm$0.43 45.17 $\pm$0.11 44.21 $\pm$0.10 48.75 $\pm$0.37
Zoom Blur 29.86 39.24 $\pm$0.26 36.62 $\pm$0.11 40.53 $\pm$0.19 40.26 $\pm$0.21 45.16 $\pm$0.08
Snow 31.64 38.91 $\pm$0.29 38.66 $\pm$0.19 42.71 $\pm$0.11 42.31 $\pm$0.18 45.74 $\pm$0.15
Frost 33.26 42.04 $\pm$0.15 40.95 $\pm$0.20 44.28 $\pm$0.14 44.83 $\pm$0.13 48.09 $\pm$0.05
Fog 24.26 31.06 $\pm$0.18 29.22 $\pm$0.24 34.78 $\pm$0.11 34.08 $\pm$0.13 41.21 $\pm$0.19
Brightness 40.48 49.26 $\pm$0.12 47.78 $\pm$0.44 51.61 $\pm$0.39 52.16 $\pm$0.08 55.75 $\pm$0.24
Contrast 1.72 3.06 $\pm$0.09 3.58 $\pm$0.07 7.24 $\pm$0.09 5.43 $\pm$0.08 13.06 $\pm$0.18
Elastic Transform 23.85 34.57 $\pm$0.05 32.11 $\pm$0.10 35.16 $\pm$0.05 35.11 $\pm$0.05 41.69 $\pm$0.12
Pixelate 22.35 37.73 $\pm$0.18 29.02 $\pm$0.22 41.11 $\pm$0.18 37.47 $\pm$0.21 45.83 $\pm$0.17
JPEG Compression 27.65 41.92 $\pm$0.45 36.51 $\pm$0.12 44.58 $\pm$0.04 42.93 $\pm$0.08 46.55 $\pm$0.26
Mean 20.92 29.78 26.96 32.90 31.66 37.69
[…]
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-S NeurIPS’24 CLIP-OT Ours
Tiny-ImageNet 71.34 74.09 $\pm$0.12 74.51 $\pm$0.20 74.57 $\pm$0.15 75.04 $\pm$0.21 75.01 $\pm$0.09
Gaussian Noise 17.26 26.60 $\pm$0.16 23.22 $\pm$0.12 29.27 $\pm$0.03 27.98 $\pm$0.14 38.14 $\pm$0.09
Shot Noise 21.57 28.93 $\pm$0.27 27.39 $\pm$0.21 33.01 $\pm$0.18 31.47 $\pm$0.04 41.02 $\pm$0.13
Impulse Noise 17.32 $2$$\pm$0.29 19.79 $\pm$0.18 27.93 $\pm$0.09 25.38 $\pm$0.15 35.48 $\pm$0.10
Defocus Blur 35.63 38.66 $\pm$0.19 39.44 $\pm$0.27 39.99 $\pm$0.15 40.55 $\pm$0.04 46.03 $\pm$0.17
Glass Blur 13.62 10.98 $\pm$0.13 15.72 $\pm$0.18 17.09 $\pm$0.15 17.44 $\pm$0.16 27.93 $\pm$0.18
Motion Blur 50.11 52.53 $\pm$0.01 53.91 $\pm$0.32 53.63 $\pm$0.28 54.46 $\pm$0.16 57.78 $\pm$0.06
Zoom Blur 43.65 46.02 $\pm$0.15 48.82 $\pm$0.11 47.39 $\pm$0.04 48.28 $\pm$0.09 52.67 $\pm$0.09
Snow 48.04 49.13 $\pm$0.33 51.50 $\pm$0.27 51.97 $\pm$0.14 52.60 $\pm$0.04 56.85 $\pm$0.13
Frost 49.36 50.46 $\pm$0.09 53.89 $\pm$0.22 51.49 $\pm$0.11 53.77 $\pm$0.16 57.56 $\pm$0.35
Fog 37.58 38.63 $\pm$0.15 40.43 $\pm$0.11 42.35 $\pm$0.04 42.57 $\pm$0.09 50.49 $\pm$0.09
Brightness 55.85 59.25 $\pm$0.04 61.07 $\pm$0.09 60.69 $\pm$0.12 62.01 $\pm$0.04 65.28 $\pm$0.31
Contrast 5.78 8.66 $\pm$0.14 7.48 $\pm$0.18 11.25 $\pm$0.08 9.21 $\pm$0.15 21.57 $\pm$0.13
Elastic Transform 34.15 36.81 $\pm$0.19 40.65 $\pm$0.27 38.44 $\pm$0.22 39.85 $\pm$0.08 47.86 $\pm$0.17
Pixelate 45.90 50.77 $\pm$0.18 52.28 $\pm$0.31 52.22 $\pm$0.06 51.83 $\pm$0.04 58.23 $\pm$0.29
JPEG Compression 50.13 51.88 $\pm$0.13 52.38 $\pm$0.24 52.46 $\pm$0.19 54.19 $\pm$0.16 56.82 $\pm$0.24
Mean 34.98 40.28 41.07 42.98 43.28 50.36
[…]
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-S NeurIPS’24 CLIP-OT Ours
Tiny-Imagenet 58.29 57.72 $\pm$0.14 58.90 $\pm$0.15 59.85 $\pm$0.01 61.35 $\pm$0.17 63.69 $\pm$0.15
Gaussian Noise 7.08 8.01 $\pm$0.05 9.29 $\pm$0.03 14.44 $\pm$0.25 13.02 $\pm$0.07 21.40 $\pm$0.37
Shot Noise 9.41 10.04 $\pm$0.07 11.70 $\pm$0.07 17.44 $\pm$0.12 15.94 $\pm$0.07 24.90 $\pm$0.17
Impulse Noise 3.44 4.18 $\pm$0.02 4.85 $\pm$0.01 10.37 $\pm$0.16 6.90 $\pm$0.01 17.34 $\pm$0.14
Defocus Blur 21.71 24.53 $\pm$0.05 27.56 $\pm$0.10 31.46 $\pm$0.26 29.91 $\pm$0.13 35.39 $\pm$0.16
Glass Blur 9.12 10.09 $\pm$0.06 11.03 $\pm$0.10 15.84 $\pm$0.09 14.01 $\pm$0.10 21.16 $\pm$0.19
Motion Blur 34.52 36.94 $\pm$0.07 38.97 $\pm$0.05 41.34 $\pm$0.20 41.26 $\pm$0.04 46.26 $\pm$0.37
Zoom Blur 27.44 29.48 $\pm$0.06 34.29 $\pm$0.07 35.06 $\pm$0.15 33.96 $\pm$0.07 40.93 $\pm$0.05
Snow 32.51 32.20 $\pm$0.03 34.45 $\pm$0.10 36.79 $\pm$0.08 37.76 $\pm$0.09 42.32 $\pm$0.14
Frost 36.33 35.72 $\pm$0.02 37.13 $\pm$0.05 38.37 $\pm$0.13 39.65 $\pm$0.03 44.97 $\pm$0.17
Fog 25.94 27.46 $\pm$0.04 28.89 $\pm$0.08 33.51 $\pm$0.19 32.13 $\pm$0.08 38.60 $\pm$0.32
Brightness 40.15 39.79 $\pm$0.08 43.31 $\pm$0.10 46.52 $\pm$0.06 46.93 $\pm$0.13 53.10 $\pm$0.05
Contrast 1.81 2.24 $\pm$0.04 3.15 $\pm$0.04 6.07 $\pm$0.12 3.53 $\pm$0.02 11.88 $\pm$0.23
Elastic Transform 30.40 31.92 $\pm$0.05 33.88 $\pm$0.14 33.74 $\pm$0.11 35.01 $\pm$0.13 40.73 $\pm$0.16
Pixelate 22.78 24.79 $\pm$0.03 27.70 $\pm$0.12 34.84 $\pm$0.09 31.55 $\pm$0.08 41.84 $\pm$0.09
JPEG Compression 29.59 30.93 $\pm$0.13 33.60 $\pm$0.15 37.29 $\pm$0.14 36.46 $\pm$0.11 42.86 $\pm$0.04
Mean 22.14 23.22 25.32 28.87 27.87 34.91
[…]